\begin{thebibliography}{99}

\bibitem{allouba2} H. Allouba, \emph{Brownian-time processes: The pde connection and the corresponding
Feynman-Kac formula,} Trans. Amer. Math. Soc. 354 (2002), no.11
4627 - 4637.

\bibitem{allouba1} H. Allouba and W. Zheng, \emph{Brownian-time processes: The pde connection and the
half-derivative generator,} Ann. Prob.  29 (2001), no. 2,
1780-1795.

%\bibitem{applebaum} D. Applebaum (2004) \emph{L\'{e}vy Processes and Stochastic Calculus.} Cambridge studies in advanced mathematics.

%\bibitem{ABHN} W. Arendt, C. Batty, M. Hieber, and F. Neubrander, {\it  Vector-valued Laplace transforms and Cauchy problems.} Monographs in Mathematics, Birkhaeuser-Verlag, Berlin (2001).

%\bibitem{Bachelier} L.J.B. Bachelier (1900) {\it Th\'eorie de la Sp\'eculation.} Gauthier-Villars, Paris.

\bibitem{Bajlekova2001} E.G. Bajlekova, \emph{Fractional evolution equations in Banach spaces}, Ph.D. thesis, Eindhoven University of Technology, 2001.

%\bibitem{Barlow90} M.T. Barlow (1990) Random walks and diffusion on fractals. In {\it Proceedings of the Internatioal Congress of Mathematicians, Kyoto, Japan} {\bf 2}, 1025--1035.  Springer-Verlag.

%\bibitem{BarlowBass} M.T. Barlow and R.F. Bass (1993) Coupling and Harnack inequalities for Sierpi\'nski carpets. {\it Bull. Amer. Math. Soc.} {\bf 29} 208--212.

%\bibitem{BarlowParkins} M.T. Barlow and E.A. Perkins (1988) Brownian motion on the Sierpi\'nski carpet.  {\it Probab. Theory Rel. Fields} {\bf 79} 543--623.

\bibitem{fracCauchy} B. Baeumer and M.M. Meerschaert,
\emph{Stochastic solutions for fractional Cauchy problems,}
Fractional Calculus Appl. Anal. 4 (2001), 481-500.

\bibitem{nane-m-b} B. Baeumer, M.M. Meerschaert and E. Nane, \emph{Brownian subordinators and fractional Cauchy problems:} Submitted (2007).

\bibitem{bandeb} R. Ba\~{n}uelos and  R.D. DeBlassie (2006), The exit distribution for iterated Brownian motion in
cones. {\it Stochastic Processes and their Applications} {\bf 116}
no. 1, 36--69.


\bibitem{coupleCTRW} P. Becker-Kern, M.M. Meerschaert and H.P.
Scheffler (2004) Limit theorems for coupled continuous time random walks.  {\it The Annals of Probability} {\bf 32}, No. 1B, 730--756.

%\bibitem{WRRapply} D. Benson, S. Wheatcraft and M. Meerschaert (2000) Application of a fractional advection-dispersion equation. {\it Water Resour. Res.} {\bf 36}, 1403--1412.

%\bibitem{WRRtheory} D. Benson, S. Wheatcraft and M. Meerschaert (2000) The fractional-order governing equation of L\'evy motion, {\it Water Resources Research} {\bf 36}, 1413--1424.

%\bibitem{BensonTiPM} D. Benson, R. Schumer, M. Meerschaert and S. Wheatcraft (2001) Fractional dispersion, L\'evy motions, and the MADE tracer tests. {\it Transport in Porous Media} {\bf 42}, 211--240.

\bibitem{bertoin} J. Bertoin (1996) \newblock {\it L\'evy processes}. Cambridge University Press.

%\bibitem{Bisquert} J. Bisquert (2003) Fractional Diffusion in the Multiple-Trapping Regime and Revision of the Equivalence with the Continuous-Time Random Walk.  {\it Physical Review Letters} {\bf 91}, No. 1, 602--605.

\bibitem{burdzy1} K. Burdzy, \emph{Some path properties of iterated Brownian
motion,} In Seminar on Stochastic Processes (E. \c{C}inlar, K.L.
Chung and M.J. Sharpe, eds.), Birkh\"{a}user, Boston, (1993),
67-87.

\bibitem{burdzy2}  K. Burdzy, \emph{Variation of iterated
Brownian motion,} In Workshops and Conference on Measure-valued
Processes, Stochastic Partial Differential Equations and
Interacting Particle Systems (D.A. Dawson, ed.)
 Amer. Math. Soc. Providence, RI, (1994),35-53.

\bibitem{burdzy-khos} K. Burdzy and D. Khoshnevisan, \emph{The level set of iterated Brownian
motion,} S\'{e}minarie de probabilit\'{e}s XXIX (Eds.: J
Az\'{e}ma, M. Emery, P.-A. Meyer and M. Yor), Lecture Notes in
Mathematics, 1613, Springer, Berlin, (1995), 231-236.

\bibitem{bukh} K. Burdzy and D. Khoshnevisan, \emph{Brownian motion in a Brownian crack,} Ann. Appl. Probabl.
8 (1998), no. 3, 708-748.

\bibitem{CCFR} E. Cs\'{a}ki, M. Cs\"{o}rg\"{o}, A. F\"{o}ldes, and
P. R\'{e}v\'{e}sz, \emph{The local time of iterated Brownian
motion,} J. Theoret. Probab. 9 (1996), 717-743.

\bibitem{deblassie}  R. D. DeBlassie, \emph{Iterated Brownian
motion in an open set,} Ann. Appl. Prob. 14 (2004), no. 3,
1529-1558.

%\bibitem{Einstein} A. Einstein (1956) {\it Investigations on the theory of Brownian movement}.  Dover, New York.

%\bibitem{FIN} L.R.G. Fontes, M. Isope and C.M. Newman (2002) Random walks with strongly inhomogeneous rates and singular diffusions:  Convergence, localization, and aging in one dimension.  {\it Ann. Probab.}
%{\bf 30}, No. 2, 579–-604.

%\bibitem{funaki}  T. Funaki, \emph{A probabilistic
 %construction of the solution of some higher order parabolic differential
 %equations,}
%Proc. Japan Acad. Ser. A. Math. Sci. 55 (1979), no. 5, 176-179.

%\bibitem{GMSR} R. Gorenflo, F. Mainardi, E. Scalas and M. Raberto (2001) Fractional calculus and continuous-time finance. III. The diffusion limit. Mathematical finance (Konstanz, 2000), 171--180, {\it Trends Math.}, Birkh\"'{a}user, Basel.

\bibitem{hawkes}J. Hawkes, \emph{A lower Lipschitz condition for stable subordinator}. Z. Wahrsch. Verw. Gebiete 17 (1971), 23-32.

%\bibitem{HiPh} E. Hille and R.S. Phillips  (1957) {\it Functional Analysis and Semi-Groups.} Amer. Math. Soc. Coll. Publ. {\bf 31}, American Mathematical Society, Providence.

\bibitem{hu-2} Y. Hu, \emph{Hausdorff and packing measures of the level sets of iterated Brownian motion.} J. Theoret. Probab. 12 (1999), no. 2, 313-346.

\bibitem{hu} Y. Hu., D. Pierre-Loti-Viaud, and Z. Shi, \emph{Laws of iterated logarithm for iterated Wiener
proceses}, J. Theoret. Probabl. 8 (1995), 303-319.

%\bibitem{Jacob} N. Jacob (1996) {\it Pseudo-Differential Operators and
%Markov Processes}. Berlin : Akad. Verl.

%\bibitem{Jacob98} N. Jacob (1998) Characteristic functions and symbols in the theory of Feller processes. {\it Potential Anal.} {\bf 8}, no. 1, 61--68.

%\bibitem{JSMarkov} N. Jacob and R. Schilling (2001) L\'evy-type processes and pseudodifferential operators. {\it L\'evy processes}, 139--168, Birkh\"auser, Boston, MA.

%\bibitem{JurekMason} Z. Jurek and J.D. Mason (1993) {\it Operator-Limit Distributions in Probability Theory}. Wiley, New York.

\bibitem{klewis} D. Khoshnevisan and T.M. Lewis, \emph{Stochastic calculus for Brownian motion in a Brownian
fracture,} Ann. Applied Probabl. 9 (1999), no. 3, 629-667.

\bibitem{koslew} D. Khoshnevisan and T.M. Lewis, \emph{Chung's law of the iterated logarithm for
iterated Brownian motion,} Ann. Inst. H. Poincar\'{e} Probab.
Statist. 32 (1996), no. 3, 349-359.

\bibitem{koch1}A. N. Kochubei, \emph{A Cauchy problem for evolution equations of fractional
order,} Differential Equations, 25 (1989),  967-974.

\bibitem{koch2}A. N. Kochubei, \emph{Fractional-order diffusion,} Differential Equations,  26
(1990), 485-492.


%\bibitem{RVbook}
%M.M. Meerschaert and H.P. Scheffler (2001)
%\newblock {\it Limit Distributions for Sums of Independent Random Vectors: Heavy Tails in Theory and Practice}.
%\newblock Wiley Interscience, New York.

\bibitem{Zsolution} M.M. Meerschaert, D.A. Benson, H.P. Scheffler and B. Baeumer (2002)  Stochastic solution of space-time fractional diffusion equations. {\it Phys. Rev. E} {\bf 65}, 1103--1106.

%\bibitem{couplePR} M.M. Meerschaert, D.A. Benson, H.P. Scheffler and P. Becker-Kern (2002) Governing equations and solutions of anomalous random walk limits. {\it Phys. Rev. E} {\bf 66}, 102R-105R.

\bibitem{limitCTRW} M.M. Meerschaert and H.P. Scheffler (2004)
Limit theorems for continuous time random walks with infinite mean
waiting times. {\it J. Applied Probab.} {\bf 41}, No. 3, 623--638.

%\bibitem{ultraslow2}  M.M. Meerschaert and H.P. Scheffler(2006) Stochastic model for ultraslow diffusion.  {\it Stoch. Proc. Appl.}, {\bf 116}, No.\ 9, 1215--1235.

%\bibitem{coupleEcon} M.M. Meerschaert and E. Scalas (2006) Coupled continuous time random walks in finance. {\it Physica A}, {\bf 370}, 114--118.

%\bibitem{MetzlerK} R. Metzler and J. Klafter (2000) The random walk's guide to anomalous diffusion: A fractional dynamics approach. {\it Phys. Rep.} {\bf 339}, 1--77.

%\bibitem{MetzlerKlafter} R. Metzler and J. Klafter (2004) The restaurant at the end of the random walk: recent developments in the description of anomalous transport by fractional dynamics.  {\it J. Physics A} {\bf 37},  R161–-R208.



\bibitem{nane} E. Nane, \emph{Iterated Brownian motion in parabola-shaped domains,} Potential Analysis, 24 (2006), 105-123.

\bibitem{nane2} E. Nane, \emph{Iterated Brownian motion in bounded domains in
$\RR{R}^{n}$,} Stochastic Processes and Their Applications, 116
(2006), 905-916.

\bibitem{nane3} E. Nane, \emph{Higher order PDE's and iterated processes},  Trans. American Math. Soc. (to appear).

\bibitem{nane4} E. Nane, \emph{Laws of the iterated logarithm for $\alpha$-time Brownian motion}, Electron. J. Probab.
11 (2006), no. 18, 434--459 (electronic).


\bibitem{nane5} E. Nane, \emph{Isoperimetric-type inequalities for iterated Brownian motion in $\RR{R}^{n}$},
Stat. Probab. Lett. (to appear) , math.PR/0602188.

\bibitem{nane6} E. Nane, \emph{ Lifetime asymptotics of iterated Brownian motion in $\RR{R}^{n}$,} Esaim:PS, March 2007, Vol. 11, pp. 147-160.

\bibitem{nigmatullin} R. R. Nigmatullin, \emph{The realization of the generalized transfer in a medium
with fractal geometry.} Phys. Status Solidi B, 133 (1986), 425-430.


\bibitem{OB}  E. Orsingher and L. Beghin (2004) Time-fractional telegraph equations and telegraph processes with Brownian time.  {\it Prob. Theory Rel. Fields} {\bf 128}, 141--160.

%\bibitem{pazy} A. Pazy  (1983) {\it Semigroups of Linear Operators andApplications to Partial Differential equations.} Applied MathematicalSciences {\bf 44}, Springer-Verlag, New York.

%\bibitem{Prosen} T. Prosen and M. Znidaric (2001) Anomalous diffusion and dynamical localization in polygonal billiards.  {\it Phys. Rev. Lett.} {\bf 87}, 114101--114104.

%\bibitem{Rudin} W. Rudin (1973) Functional Analysis.  2nd Edition, McGraw-Hill, New York.

%\bibitem{sam-taqqu} G. Samorodnitsky and M. Taqqu, \emph{Stable non-Gaussian Random processes,} Chapman and Hall, New york (1994).


%\bibitem{sato}
%K.I. Sato (1999)
%\newblock {\it L\'evy Processes and Infinitely Divisible Distributions}.
%\newblock Cambridge University Press.

%\bibitem{scalas1} E. Scalas (2004) Five years of Continuous-Time Random Walks in Econophysics.{\it Proceedings of WEHIA 2004}, A. Namatame (ed.), Kyoto.


%\bibitem{schilling}Schilling, R.L.: Growth and H\"older conditions for sample paths of Feller proceses.\newblock Probability Theory and Related Fields \textbf{112}, 565--611 (1998)

%\bibitem{Eulerian} R. Schumer, D.A. Benson, M.M. Meerschaert and S. W. Wheatcraft (2001) Eulerian derivation of the fractional advection-dispersion equation. {\it J. Contaminant Hydrol.}, {\bf 48}, 69--88.

\bibitem{shi-yor} Z. Shi and M. Yor, \emph{Integrability and lower limits of the local time of iterated Brownian motion,} Studia Sci. Math. Hungar. 33 (1997), no. 1-3, 279-298.

\bibitem{sch-wyss}W. R. Schneider and W. Wyss, \emph{Fractional diffusion and wave equations,} J.
Math. Phys., 30 (1989), 134-144.


%\bibitem{SKW} M. Shlesinger, J. Klafter and Y.M. Wong (1982) Random walks with infinite spatial and temporal moments. {\it J. Statist. Phys.} {\bf 27}, 499--512.

%\bibitem{Sinai} Y.G. Sinai (1982)  The limiting behavior of a one-dimensional randomwalk in a random medium.  {\it Theor. Probab. Appl.} {\bf 27}, 256--268.

%\bibitem{SokolovKlafter} I.M. Sokolov and J. Klafter (2005) From Diffusion to Anomalous Diffusion: A Century after Einstein's Brownian Motion. {\it Chaos} {\bf 15}, 6103--6109.

\bibitem{xiao} Y. Xiao, \emph{Local times and related properties of multi-dimensional iterated Brownian
motion,} J. Theoret. Probab. 11 (1998), no. 2, 383-408.

\bibitem{zaslavsky} G. Zaslavsky, \emph{Fractional kinetic equation for Hamiltonian chaos. Chaotic advection,
tracer dynamics and turbulent dispersion.} Phys. D 76 (1994),
110-122.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath, amsthm}
\usepackage{amsmath,amssymb,amsthm}
\usepackage{amsfonts}
\usepackage[mathscr]{eucal}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem]{Corollary}

\theoremstyle{definition}

\theoremstyle{remark}

\numberwithin{equation}{section}

\newcommand{\RR}[1]{\mathbb{#1}}

\newcommand{\rd}{{\mathbb R^d}}

\begin{document}

\title{\bf Symmetric $\alpha$-stable subordinators and Cauchy problems}

\author{Erkan Nane}
\address{Erkan Nane, Department Statistics and Probability,
Michigan State University, East Lansing, MI 48823}
\email{nane@stt.msu.edu}
\urladdr{http://www.stt.msu.edu/$\sim$nane}


\begin{abstract}
 
 We survey the results in Nane (E. Nane, \emph{Higher order PDE's and iterated processes},  Trans. American Math. Soc. (to appear)) and Baeumer, Meerschaert, and Nane (B. Baeumer, M.M. Meerschaert and E. Nane, \emph{Brownian subordinators and fractional Cauchy problems:} Submitted (2007)) which deal with PDE connection of some iterated processes, and obtain a new probabilistic proof of the equivalence of the higher order PDE's   and fractional in time PDE's.
\end{abstract}

\keywords{Iterated Brownian motion, PDE connection,
$\alpha-$ stable process, $\alpha$-time process,  L\'{e}vy process, Brownian subordinator, Caputo derivative, fractional derivative in time}

\subjclass{60J65, 60K99.}
\maketitle

\section{\bf Introduction }

In recent years, starting with the articles of Burdzy
\cite{burdzy1, burdzy2}, researchers have shown interest in iterated
processes in which one changes the time parameter with
one-dimensional Brownian motion.

To define {\bf iterated  Brownian motion} $Z_{t}$, due to
Burdzy \cite{burdzy1}, started at $z \in \RR{R}$,
 let $X_{t}^{+}$, $X_{t}^{-}$ and $Y_{t}$
 be three  independent
one-dimensional Brownian motions, all started at $0$. {\bf Two-sided
Brownian motion} is defined to be
\[ X_{t}=\left\{ \begin{array}{ll}
X_{t}^{+}, &t\geq 0\\
X_{(-t)}^{-}, &t<0.
\end{array}
\right. \] Then iterated Brownian motion started at $z \in \RR{R}$
is
$ Z_{t}=z+X(Y_{t}),\ \ \    t\geq 0.$


\subsection{\bf BM versus IBM}
This process is not Markovian or Gaussian but, it has many properties analogous to those of
Brownian motion. We list a few;

 IBM $Z_{t}$ has stationary (but not independent) increments, and is a
{\bf self-similar process} of index $1/4$.
{\bf Laws of the iterated logarithm (LIL)} holds:
usual LIL by Burdzy \cite{burdzy1} shows
$$
\limsup_{t\to\infty}\frac{Z(t)}{t^{1/4}(\log \log
(1/t))^{3/4}}=\frac{2^{5/4}}{3^{3/4}} \ \ \  a.s.
$$
 Chung-type LIL is obtained by Khoshnevisan and Lewis \cite{koslew} and Hu et al.
\cite{hu}.
 Khoshnevisan
and Lewis \cite{klewis} extended results of Burdzy \cite{burdzy2}, to develop a
{\bf stochastic calculus} for iterated Brownian motion.
 Burdzy and Khosnevisan \cite{bukh} showed that IBM can be used to model diffusion in a crack.
 Local times of this process was  studied by Burdzy
and Khosnevisan \cite{burdzy-khos},  Cs\'{a}ki, Cs\"{o}rg\"{o}, F\"{o}ldes,
and R\'{e}v\'{e}sz \cite{CCFR}, Shi and Yor \cite{shi-yor}, Xiao \cite{xiao}, and Hu \cite{hu-2}.
 Ba\~{n}uelos and DeBlassie \cite{bandeb} studied the
{\bf distribution of exit place} for iterated Brownian motion
in cones.
 DeBlassie \cite{deblassie} studied the lifetime asymptotics of iterated
Brownian motion in cones and Bounded domains. Nane \cite{nane, nane2, nane5, nane6}
extended some of the results of DeBlassie. 

\subsection{\bf  PDE connection}

In addition to the above properties of IBM, there is an interesting
connection between iterated Brownian motion
 and the {\bf biharmonic operator $\Delta ^{2}$}. Allouba and Zheng \cite{allouba1} show that if we replace the
outer process $X(t)$ in the definition of iterated Brownian motion  with a continuous Markov process with  $L_x$ as the semigroup generator, then
$u(t,x)=E_{x}[f(Z_{t})]:=E[f(Z_{t})|Z_0=x]$
solves the Cauchy initial value problem
\begin{equation}\label{ibm-pde00}
\frac{\partial}{\partial t}u(t,x) =
\frac{{L_x}f(x)}{\sqrt{\pi t}}+ {L_x}^{2}u(t,x); \quad u(0,x) = f(x), \ t>0,  x\in \rd .
\end{equation}
When $Z_t$ is an iterated Brownian motion, this was also obtained by DeBlassie \cite{deblassie} by a different method.
Let $Z^1_t=X(|Y_t|)$, then Allouba and Zheng \cite{allouba2} shows $E_{x}[f(Z_{t})]=E_{x}[f(Z^1_{t})]$.

\subsection{\bf Fractional Cauchy problems}

Nigmatullin \cite{nigmatullin} gave a physical derivation of the fractional kinetic
equation for some special $\beta$
\begin{equation}\label{frac-derivative-0}
\frac{\partial^\beta}{\partial t^\beta}u(t,x) =
{L_x}u(t,x); \quad u(0,x) = f(x)
\end{equation}
where $0< \beta <1$ and $L_x$ is the generator of
some continuous Markov process $X_0(t)$ started at $x=0$. Here $\partial^{\beta} g(t)/\partial t^\beta $
is the Caputo fractional derivative in time, which can be defined as the
inverse Laplace transform of $s^{\beta}\tilde{g}(s)-s^{\beta -1}g(0)$, with
$\tilde{g}(s)=\int_{0}^{\infty}e^{-st}g(t)dt$ the usual Laplace transform.

Mathematical study of equation (\ref{frac-derivative-0}) was initiated by \cite{sch-wyss, koch1, koch2}. The existence and uniqueness of solutions to equation (\ref{frac-derivative-0}) was proved in \cite{koch1, koch2}.
This equation was also used by Zaslavsky \cite{zaslavsky} for Hamiltonian chaos.


Baeumer and Meerschaert \cite{fracCauchy} and Meerschaert and Scheffler \cite{limitCTRW}
 show that the fractional Cauchy problem (\ref{frac-derivative-0}) is related to a certain
 class of subordinated stochastic processes;
 take $D_t$ to be the stable subordinator,
 a L\'evy process with strictly increasing sample paths such that $E[e^{-sD_t}]=e^{-ts^\beta}$,
  see for example Bertoin \cite{bertoin}.  Define the inverse or hitting time or first passage time process
\begin{equation}\label{Edef}
E_t=\inf\{x>0: D(x)>t\} .
\end{equation}
The subordinated process $Z_t=X_0(E_{t})$ occurs as the scaling limit of a
 continuous time random walk (also called a renewal reward process), in which iid
 random jumps are separated by iid positive waiting times (Meerschaert and Scheffler (2004)\cite{limitCTRW}).
Theorem 3.1 in Baeumer and Meerschaert \cite{fracCauchy} shows that, in the
case $p(t,x)=T(t)f(x)$ is a bounded continuous semigroup on a
Banach space, the formula
$$
u(t,x)=\int_0^\infty p((t/s)^\beta,x) g_\beta(s)\,ds
=
\frac{t}{\beta}\int_{0}^{\infty}p(x,s)g_{\beta}(\frac{t}{s^{1/\beta}})s^{-1/\beta
-1}ds \nonumber 
$$
yields a solution to the   fractional Cauchy problem (\ref{frac-derivative-0}).
  Here, $g_\beta(t)$ is the smooth
density of the stable subordinator, with $\tilde g_\beta(s)=\int_0^\infty
e^{-st}g_\beta(t)\,dt=e^{-s^\beta}$.

\section{\bf Brownian subordiantors and  fractional Cauchy problems}
We give a probabilistic proof of the following theorem. A variation of this result was realized by Orsingher and Benghin \cite{OB} for a version of iterated Brownian motion.
\begin{theorem}[Baeumer, Meerschaert and Nane (2007)\cite{nane-m-b}]\label{equiv-pde}
Let $L_{x}$ be the generator of a Markov semigroup $T(t)f(x)=E_{x}[f(X_{t})]$,
and take $f\in D(L_{x})$ the domain of the generator.  Then, both the higher order Cauchy problem (\ref{ibm-pde00})
 and the fractional Cauchy problem (\ref{frac-derivative-0})
 with $\beta=1/2$,
have the same solution
\begin{eqnarray}
u(t,x)& = &E_x[f(Z_t)]
= \frac{2}{\sqrt{4\pi t}}\int_{0}^{\infty}T(s)f(x)\exp\left(-\frac{s^2}{4t}\right)ds. \label{ACPsoln1}
\end{eqnarray}
\end{theorem}
\begin{proof}
$E_t$ is the inverse of a $1-1/\alpha$ stable subordinator. $E_t$ then is the local time of symmetric stable process of index $\alpha$. In the case $\alpha=2$, local time of Brownian motion is the same as $\sup_{0<s<t} B_s$. On the other hand, $\sup_{0<s<t} B_s$ and $|B_t|$  are same in distribution by the reflection principle. Hence, $E_t$ and $|B_t|$ have the same one-dimensional distributions, implying the result of the theorem.
\end{proof}

We obtain the following corollary of our theorem

\begin{corollary}[Baeumer, Meerschaert and Nane (2007)\cite{nane-m-b}] 
For any continuous Markov process
$X(t)$, both the Brownian-time subordinated process $X(|Y_t|)$ and
the process $X(E_t)$ subordinated to the inverse $1/2$-stable
subordinator have the same one-dimensional distributions.  Hence
they are both stochastic solutions to the fractional Cauchy
problem (\ref{frac-derivative-0}), or equivalently, to the higher
order Cauchy problem \eqref{ibm-pde00}.
\end{corollary}
In contrast to the previous corollary, we have
\begin{theorem}
Let $Y$ be a symmetric stable process of index $1<\alpha <2$, and $E_t$ is the inverse of a stable subordinator of index $1-1/\alpha$.
The processes $X(E_t)$ and $X(|Y_t|)$ do not have same one-dimensional distribution.
\end{theorem}
\begin{proof}Let $L_1^0$ be the local time at $x=0$. $L_1^0$ has the same one-dimenional distributions as $E_t$.
Lemma 1 in Hawkes \cite{hawkes} implies that
\begin{equation}\label{local-tail}
P[L_1^0>\lambda]\sim C_1\lambda^{-\alpha/2}\exp(-C_{\alpha h}\lambda^{\alpha}).
\end{equation}
 Proposition 4 in Bertoin \cite{bertoin} shows
\begin{equation}\label{stable-tail}
 P[Y_1>u]\sim P[\sup_{0\leq s \leq 1}Y_s >u]\sim cu^{-\alpha}.
\end{equation}
 
 The results in equations (\ref{local-tail}) and (\ref{stable-tail})establish  that in the case $Y_t$ is a symmetric stable process of index $\alpha <2$, $|Y_t|$ and $E_t$ do not have the same one-dimensional distributions.
\end{proof}

When the outer proces is   L\'evy process we have uniqueness of the solutions in Theorem \ref{equiv-pde}.
 The proof  relies on a Laplace-Fourier transform argument.

\begin{theorem}[Baeumer, Meerschaert and Nane (2007)\cite{nane-m-b}]\label{unique-sol}
Suppose that $X(t)=x+X_0(t)$ where $X_0(t)$ is a L\'evy process starting at zero.  If $L_x$ is the generator of the semigroup $T(t)f(x)=E_x[(f(X_t))]$ on $L^1(\rd)$, then for any $f\in D(L_{x})$, both the initial value problem \eqref{ibm-pde00}, and the fractional Cauchy problem \eqref{frac-derivative-0} with $\beta=1/2$, have the same unique solution given by \eqref{ACPsoln1}.
\end{theorem}

An easy extension of the argument for Theorem \ref{unique-sol} shows that, under the same conditions, for any $n=2,3,4,\ldots$ both the Cauchy problem
\begin{equation}\begin{split}\label{one-third-n}
\frac{\partial u(t,x)}{\partial t} & =\sum_{j=1}^{n-1} \frac{t^{1-j/n}}{\Gamma(j/n)} L_x^j f(x)
+ L_x^n u(t,x); u(0,x)=  f(x)\quad 
\end{split}\end{equation}
and the fractional Cauchy problem (\ref{frac-derivative-0}) with $\beta=1/n$
have the same unique solution given by
$
u(t,x) =\int_0^\infty p((t/s)^\beta,x) g_\beta(s)\,ds
$
with $\beta=1/n$.  
Hence the process $Z_t=X(E_t)$ is also the stochastic solution to this higher order Cauchy problem.


\section{\bf  Other subordinators}

An $\alpha$-time process is a Markov process subordinated to the absolute value of an independent one-dimensional symmetric
$\alpha$-stable process: $Z_t=B(|S_t|)$, where $B_t$ is a Markov process and $S_t$ is an independent symmetric $\alpha$-stable process both started at $0$. Let $Z_t^x =x+ Z_t$ the process started at $x$.

This process is self similar with index $1/2\alpha$ when the outer
process $X$ is a Brownian motion. In this case, Nane \cite{nane4} defined
the local time of this process and obtained laws of the iterated
logarithm for the local time for large time.

\subsection{\bf PDE-connection:}

\begin{theorem}[Nane (2005) \cite{nane3}]

Let $T(s)f(x)=E[f(X^{x}(s))]$ be the semigroup of the
continuous Markov process $X^{x}(t) $ and let $ L_x$ be its
generator. Let $\alpha=1$. Let $f$ be a bounded measurable
function in the domain of $ L_x$, with $D_{ij}f$ bounded and
H\"{o}lder continuous for all $1\leq i,\ j \leq n$. Then
$u(t,x)=E[f(Z^{x}_t)]$
solves
\begin{eqnarray}
\frac{\partial ^{2}}{\partial t^{2}}u(t,x)\ &\ = \ & -\frac{2L_x f(x)}{\pi t}\ - \ L_x^{2}u(t,x); \ u(0,x)  =\   f(x).\nonumber
\end{eqnarray}
\end{theorem}

 For  $\alpha=l/m \neq 1$ rational: the PDE is more
complicated since kernels of symmetric $\alpha$-stable processes
satisfy a higher order PDE:
$$
(\frac{\partial^{2}}{\partial
s^{2}})^{l}+(-1)^{l+1}\frac{\partial^{2m}}{\partial
t^{2m}})p_{t}^{\alpha}(0,s)=0.
$$
We also have to assume that  we can take the operator out of the integral. This is
valid for $\alpha=1/m$, $m=2,3,\cdots$, by a Lemma
in Nane \cite{nane3}.

\begin{theorem}[Nane (2005)\cite{nane3}]
Let $\alpha \in (0,2)$ be a rational with $\alpha=l/m$, where $l$ and $m$
are relatively prime. Let $ T(s)f(x)=E[f(X^{x}(s))]$ be the
semigroup of the continuous Markov process $X^{x}(t) $ and let
$L_x$ be its generator. Let $f$ be a bounded measurable
function in the domain of $L_x$, with $D^{\gamma}f$  bounded and
H\"{o}lder continuous for all multi index $\gamma$ such that $|\gamma |=2l$. Then
$u(t,x)=E[f(Z^{x}_{t})]$
solves
\begin{eqnarray}
 (-1)^{l+1}\frac{\partial ^{2m}}{\partial t^{2m}}u(t,x)&=&
-2\sum_{i=1}^{l}\left(
\frac{\partial^{2l-2i}}{\partial s^{2l-2i}}p_{t}^{\alpha}(0,s)|_{s=0}\right)L_x^{2i-1}f(x)\ 
  -\ L_x^{2l}u(t,x); \nonumber\\
u(0,x)&   =  & f(x).\nonumber
\end{eqnarray}
\end{theorem}
For some other connections of PDE's and iterated processes, see papers by Nane \cite{nane3} and Allouba and Zheng \cite{allouba1}, Allouba \cite{allouba2}, Baeumer et al. \cite{nane-m-b} and references therein.

\section{\bf  Open Problems }

\noindent {\bf Question 1.} Looking at the governing PDE for
subordinators other than Brownian motion, are there any fractional
in time PDE which has the same solution as the higher order PDE?

\noindent {\bf Question 2.} Are there PDE connections of the
iterated processes in bounded domain as the PDE connection of
Brownian motion in bounded domains?

\section{\bf Acknowledgements:} Author thanks Professor Mark M. Meerschaert and Professor Yimin Xiao for their help and discussions on the results in this paper. I also would like to thank Professor Anatoly N. Kochubei for providing the references for the initial apperance of  equation (\ref{frac-derivative-0}) in the literature.

\end{document}